Two bars of identical mass are at 23°C. One is made from glass and the other from another substance. The specific heat capacity of glass is 840 J/(kg · C°). When identical amounts of heat are supplied to each, the glass bar reaches a temperature of 81°C, while the other bar reaches 257.5°C. What is the specific heat capacity of the other substance?
Answer #1

Given: Initial temperature for both the bars T0 = 230C ; Final temperature for glass Tg = 810C ; Final temperature for other bar T1 = 257.50C; specific heat capacity of glass Cg = 840 J/Kg0C

Now let the mass of both the bars be 'm' , specific heat capacity of of the other bar be 'C1'  and the amount of heat supplied to each of the bar be 'Q'

Then Q = m*Cg*(Tg-T0) = m*C1*(T1-T0)

or m*840*(81-23) = m*C1*(257.5-23) or C1 = 48720/234.5 = 207.76 J/Kg-0C